OpenStudy (anonymous):

Write 1.375*10^6 in standard form.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

Google it.

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

nothing came up

OpenStudy (anonymous):

Uh-huh. This is what happens when you google it: https://www.google.com/search?q=1.375*10%5E6&gws_rd=ssl

OpenStudy (anonymous):

oh I typed it different than that

OpenStudy (anonymous):

thanks

OpenStudy (igreen):

@AnswerMyQuestions You don't have Google when you take tests -.-

OpenStudy (anonymous):

...

OpenStudy (anonymous):

I didn't say you had to...

OpenStudy (igreen):

\(1.375 \times 10^6\) First we simplify \(10^6\)..which is the same thing as \(10 \times 10 \times 10 \times 10 \times 10 \times 10\) which is the same thing as \(1000000\). So we have: \(1.375 \times 1,000,000\) Now all you have to do is multiply that.. @shifuyanli

OpenStudy (igreen):

@AnswerMyQuestions If he uses Google NOW to solve his problems, he won't know how to do it when he does his test, when he does NOT have Google.